# Corner functions from entanglement indices of harmonic lattices

### v1

@@ 2 Models of harmonic lattices and entanglement indices computation
$\begin{split}\tilde{b}(3\pi/4)=0.0052,&\ \ \ \ b(3\pi/4)_{\text{std}}=0.0012,\\ \tilde{b}(3\pi/4)=0.0038,&\ \ \ \ b(3\pi/4)_{\text{std}}=0.0013.\end{split}$ (38)
The values of these corner functions are equal to the values obtained from LNs, equations (35) and (36), within a range of a few times the respective STDs. However, the ratios of $\tilde{b}(3\pi/4)$ to the values obtained from LN are 0.86 or 0.6, which is not a sufficient agreement, and as the relative standard deviations are as large as 0.23 or 0.34, they are not stable values.
At the end of the calculation for a lattice satisfying PBCs, we take a square region $D$ whose centroid is $(99.5,99.5)$ with the site at the lower left corner of $U$ as the origin and repeat the same calculations as above. The results showed that for all corner functions treated, the values and their STDs are in agreement with (35), (36), (37) and (38) up to the most significant digit of each STD that are to be expected with the universal system $U$ under PBCs.
3 Harmonic lattice with fixed end boundary conditions
In this section, as an universal system $U$, we consider a set of sites in a square satisfying FBCs of the same scale as in the previous section. Although the square has 1202 sites on a side, due to the FBCs for the two sites at each end, the number of sites with dynamic variables is 1200 and $L_{1},L_{2}=1201$. Also, for the subsystems $A_{1}$, $_{2}, A=A$$A_{1}+A_{2}$, we have the same shapes and scales as in the previous section (for $A_{2}$, square $A_{s}$, right isosceles triangle $A_{r}$, trapezoid $A_{t}$, parallelogram $A_{p}$). On the other hand, the region $D$, which includes $A$ and for which the correlation functions of Eqs.(19) and (20) are to be calculated is taken at several locations from the centroid of $U$ to near the lower left corner of $U$. We adopt the six positions of $D$ in Figure 2 as well as two regions whose centroid are (55.5,55.5) and (75.5,75.5) with the origin at the fixed end site at the lower left corner of $U$. The last two positions of $D$ are added to observe the fixed end effect to corner functions. For each of them, we compute $b_{\text{total}}$s and $\tilde{b}_{\text{total}}$s in Eqs.(31),(32) with figures of $A_{2}$ and corner functions involved as in the previous section.
@@ 3 Harmonic lattice with fixed end boundary conditions
In addition to the corner function data of the universal system obeying FBCs, we display, in the column PBC600, the values of the corner functions and their STDs of Eqs from (35) to (38) obtained from LNs and MIs corresponding to the universal system with PBCs in the previous section. These values, respectively, are almost the same as those in the column to the left 600, i.e., the values obtained in the universal system with FBCs and the $D$ region with the centroid (600.5,600.5). In each graph of Figure 3, each corner function in the column of PBC600 is also plotted at the horizontal axis value of 600, but it is indistinguishable from the plotted value of the column of FBC600, because they almost overlap. The results are expected, since, even when the universal system $U$ follows the FBCs, the effects of the fixed ends are expected to be small if the $D$ is sufficiently far from the boundary of the $U$. Furthermore, corner functions in the column of 300 to 600 (of FBC) and plotted on each line in Figure 3 have overlapping values within their STDs. Thus, for a sufficiently large system $U$, the effects of the fixed edge are to decrease in the region away from the edge of the $U$ about three to four times the size of the largest subsystem $A$, and the physical states of subsystems are expected to be the same as those of the harmonic lattice in the universal system obeying the PBCs.
position FBC55 75 100 200 300 400 500 600 PBC600
$b(\pi/2)$ 0.02966 0.02954 0.02858 0.02933 0.02948 0.02952 0.02954 0.02955 0.02954
$b(\pi/2)_{\text{std}}$ 0.00012 0.00045 0.00017 0.00018 0.00013 0.00012 0.00011 0.00011 0.00013
$b(\pi/4)$ 0.09675 0.0968 0.09546 0.09639 0.09660 0.09666 0.09668 0.09670 0.09670
$b(\pi/4)_{\text{std}}$ 0.00037 0.0014 0.00037 0.00024 0.00029 0.00031 0.00032 0.00032 0.00032
$b(3\pi/4)$ from $A_{t}$ 0.01048 0.0029 0.00627 0.00633 0.00615 0.00609 0.00606 0.00604 0.00604
$b(3\pi/4)_{\text{std}}$ 0.00099 0.0020 0.00051 0.00041 0.00044 0.00046 0.00047 0.00047 0.00047
$b(3\pi/4)$ from $A_{p}$ 0.00691 0.0052 0.00631 0.00644 0.00639 0.00637 0.00636 0.00635 0.00635
$b(3\pi/4)_{\text{std}}$ 0.00045 0.0015 0.00044 0.00032 0.00035 0.00037 0.00038 0.00038 0.00038
Table 1: Corner function values obtained from LNs at various positions in the lattice $U$.
position FBC55 75 100 200 300 400 500 600 PBC600
$\tilde{b}(\pi/2)$ 0.02986 0.03210 0.03129 0.03019 0.02987 0.02974 0.02967 0.02963 0.02962
$\tilde{b}(\pi/2)_{\text{std}}$ 0.00006 0.00024 0.00015 0.00015 0.00010 0.00007 0.00004 0.00003 0.00003
$\tilde{b}(\pi/4)$ 0.09942 0.10237 0.10194 0.10102 0.10068 0.10055 0.10046 0.10042 0.10041
$\tilde{b}(\pi/4)_{\text{std}}$ 0.00083 0.00082 0.00098 0.00094 0.00099 0.00104 0.00106 0.00109 0.00109
$\tilde{b}(3\pi/4)$ from $A_{t}$ 0.00860 0.00910 0.00774 0.00594 0.00553 0.00533 0.00527 0.00520 0.00520
$b(3\pi/4)_{\text{std}}$ 0.00138 0.00103 0.00102 0.00099 0.00108 0.00116 0.00125 0.00127 0.00124
$\tilde{b}(3\pi/4)$ from $A_{p}$ 0.00475 0.00626 0.00539 0.00423 0.00398 0.00387 0.00383 0.00377 0.00377
$\tilde{b}(3\pi/4)_{\text{std}}$ 0.00096 0.00090 0.00011 0.00011 0.00115 0.00121 0.00127 0.00127 0.00125
Table 2: Corner function values obtained from MIs at various positions in the lattice $U$.
Figure 3: Values of corner functions obtained in various $D$ regions. A scale value of the horizontal axis of each graph indicates (the $x$-coordinate of the centroid an $D$ region)0.5-0.5. The graphs in the lower row are values of $b(3\pi/4)$ obtained using $A_{t}$ and $A_{p}$, respectively.
In the conformal theories(CFTs), i.e.,the continuous limit of the above lattice theories, each corner function $\tilde{b}(\theta)$ by Renyi entropy with the order $1/2$ has a universal lower limit, $\tilde{b}\geq\frac{1}{32\pi}(\pi-\theta)^{2}$[9]. Specifically,
Looking at the values in the FBC600 columns of tables 1 and 2 and plots for those with the horizontal axis coordinate 600 of each graph in Figure 3, the error bars with STDs of $b(\pi/2)$ and $\tilde{b}(\pi/2)$ and the $b(3\pi/4)$ and $\tilde{b}(3\pi/4)$ calculated from trapezoidal $A_{2}$ in shape, $A_{t}$, overlap respectively, where the lower limits expected by the CFT are included. On the other hand, for corner functions of $\pi/4$ and those of $3\pi/4$ calculated from parallelograms, $A_{p}$, error bars of twice the standard deviation, where includes the lower limit expected by the CFT.
4 3D lattice systems: Edge contributions to entanglement
So far, taking the universal system $U$ and the subsystems $A=A_{1}+A_{2}$ on a 2D lattice, we have analyzed the LNs between A1 and A2 and the MIs. In this section, we investigate the entanglement between two adjacent sets of harmonic oscillators placed on lattice sites in $d_{U}=3$ dimensional space. For solid figures, the number of sites in the interior increases rapidly in proportion to the cube of the scale. Considering the limitation of available computer resources, we assume that the universal system $U$ is the set of sites in a cube with 120 sites on a edge. As a subsystem, we take the set of sites near the center of gravity of the system $U$ that are inside a cube or a triangular prism whose base is a isosceles right triangle. Then, assuming $U=A+B$, we compute the Renyi entropy of order $1/2$ while varying the scale (the number of sites on the shortest edges of shapes of the subsystems from 6 to 20. In these cases, the boundary surfaces of the two subsystems $A$ and $B$ have not only sharp vertices but also edges. For the entanglement entropy in the case of sharp vertices and edges at the boundary, the following equation is said to be valid from formula (6) omitting cut-off[17]:
$S_{A}^{(n)}=\alpha P_{B}+\beta_{total}\log P_{B}-\sum_{j}e(\theta_{j})L_{j}+\cdots,$ (39)
where $\beta_{total}$ is the sum of corner function $\beta(\Omega_{k})$ of sharp vertex with the solid angle $\Omega_{k}$, and $e(\theta_{j})$ and $L_{j}$ are the edge term coefficient and the length of the edges with dihedral angle $\theta_{j}$. By fitting the above data of the Renyi entropy according to the formula (39), we obtain the corner functions $\beta(\pi/2)$, $\beta(\pi/4)$, the edge term coefficients $e(\pi/2)$ and $e(\pi/4)$ and their STDs. When doing so, as is done in equations (31), (32) and later, terms proportional to $P_{B}^{-k}(k=0,1,2...,n)$ are added to the right-hand side of the equation (39), and the operations are repeated while increasing $n$. Then, we select the $n$ that gives the most stable total values of corner functions, i.e., the smallest STD of the total values. In fact, in both the PBC and FBC cases and for both $A$ shapes of cube and triangle prism, the most stable total corner functions are obtained by the fitting up to the constant term $c_{0}$.
To begin, for the case where the shape of the subsystem $A$ is a cube, let us compare the sum of the corner functions and their STDs obtained by the method described above for $S_{A}^{(1/2)}$ with those obtained by fitting to the Eq.(39) minus the edge terms. The table 3 gives the corner functions and their STDs when computing with and without edge terms for the case where $U$ satisfies the PBCs and FBCs, respectively. In the case where the edge term is considered, the coefficient values and their STDs are also included.
$\beta(\pi/2)$ $\beta(\pi/2)_{std}$ $e(\pi/2)$ $e(\pi/2)_{std}$
PBC without $e$ 0.242 0.026
with $e$ 0.005673 0.00010 0.06725 0.00003
FBC without $e$ 0.242 0.026
with $e$ 0.005604 0.000059 0.06723 0.00002
The corresponding values in the PBC row and FBC row of the table are almost the same and the corner functions without considering the edge term are about 43 times larger than those with the edge terms. The relative STDs are about 0.11 without considering the edge terms, while it is about 0.02 for PBCs and 0.01 for FBCs when the edge terms are considered. In other words, when the edge terms are taken into account, the corner functions are five to ten times more stable than when the edge terms are not taken into account. For the edge term coefficients, very stable values 0.06723, 0.06725 with relative STDs of 0.004, 0.003 are obtained for the two lattice boundary conditions. The results indicate the existence of the edge term in the Eq.(39).
Next, let us discuss the case where the shape of $A$ system is a triangular prism. This triangular prism is with the base that is an isosceles right triangle formed by dividing one face (square) of the cube described above by its diagonal, and the height being the length of one edge of the cube. This shape has four and two sharp vertices with solid angle $\pi/4$, $\pi/2$, respectively, and two and seven edges with dihedral angle $\pi/4$, $\pi/2$, respectively. For the system of this configuration, we follow the same process as before, giving the sum of the related corner functions and edge terms, and using values of $\beta(\pi/2)$ and $e(\pi/2)$ and their STDs in table 3, we obtain corner functions $\beta(\pi/4)$ and edge term coefficients $e(\pi/4)$ and their STDs as shown in table 4.
$\beta(\pi/4)$ $\beta(\pi/4)_{std}$ $e(\pi/4)$ $e(\pi/4)_{std}$
PBC 0.0115 0.0002 0.1416 0.0001
FBC 0.0114, 0.0001 0.1417 0.0001
The values of corner functions, the edge term coefficients are consistent within the range of their STDs between two cases where the universal lattice satisfies PBCs and FBCs, and both of them are stable values with relative STDs below 0.02 or 0.0008. In addition, the values of $b(\pi/4)$ and $e(\pi/4)$ are about twice as large as those of $b(\pi/2)$ and $e(\pi/2)$, respectively. This characteristics is consistent with the fact that the geometric discontinuity of the boundary surface becomes stronger as the solid angle or dihedral angle approaches an acute angle, and the contribution from sharp corners and edges increases, as was the case with the two dimensional lattice.
5 conclusions
We have studied the logarithmic negativity(LN) and mutual information(MI) between two adjacent subsets $A_{1}$ and $A_{2}$ in a isolated universal set $U$ of harmonic oscillators on a two dimensional lattice, whose continuum limit is the massless scalar field. We have let the sites of $A_{2}$ be enclosed by sites of $A_{1}$ and let the boundary of $A_{2}$ be one of various shapes with corners whose angles are $\pi/4$, $\pi/2$, $3\pi/4$. Varying the scales of subsystems $A=A_{1}+A_{2}$ and $A_{2}$ kept as similar as possible in their shapes, we obtained LNs, MIs and from them so-called corner functions for the above corner angles to verify the results of the previous study which treats corner functions mainly from LNs for the case where the canonical variables obey periodic boundary conditions(PBCs). In the case that canonical variables on the lattice obey PBCs, all the values of corner functions $b(\theta)$ obtained from LNs have almost agreed with those in the previous studies, although some of the differences are quite large relative to the standard deviations. As for the corner function values $\tilde{b}(\theta)$ obtained from MIs, the value of the corner functions $\tilde{b}(\pi/2)$ and $\tilde{b}(\pi/4)$ are almost agreed with those obtained from data of $LN$s. On the other hand, as for the values of $\tilde{b}(3\pi/4)$, considering the magnitude of the STDs and their ratio to those obtained from $LN$s data, or the previous study[11], they are not stable and enough consistent to those of the study.
[…]
@@
Furthermore, We computed the Renyi entropies of order 1/2 for a three-dimensional lattice, considering, as a subsystem, a set of sites inside a cube varying the scale of the subsystem. By fitting the data with equation (39), we obtained the corner functions and edge term coefficients as shown in table 3. We compare the STDs of the corner functions with those obtained by fitting with (39) minus the edge terms . The result shows that the STDs of the latter are 5 to 10 times larger than that of the former. This fact is the evidence of the existence of the edge term in the expansions of Renyi entropy between two adjacent systems whose boundary is a three-dimensional polyhedron. In addition, the Renyi entropy was also obtained for the set of sites in triangular prisms with vertices and edges having solid angles and dihedral angles of $\pi/4$ as the subsystem $A$. Using the results of table 3, the corner function $\beta(\pi/4)$, edge term coefficient $e(\pi/4)$ and their STDs are obtained as shown in table 4.
@@
As in the case of a two-dimensional lattice, the values of the corner functions and edge term coefficients in tables 3 and 4, where U satisfies FBCs and the $A$ system is sufficiently distant from the edge of the universal system $U$, are almost identical to the values when $U$ satisfies PBCs. In addition, these values satisfy the stability and monotonically decreasing properties with respect to the solid angle and dihedral angle, respectively, that the corner functions and edge terms coefficients would have.

### v2

@@ 2 Models of harmonic lattices and entanglement indices computation
$\begin{split}\tilde{b}(3\pi/4)=0.0052,&\ \ \ \ b(3\pi/4)_{\text{sd}}=0.0012,\\ \tilde{b}(3\pi/4)=0.0038,&\ \ \ \ b(3\pi/4)_{\text{sd}}=0.0013.\end{split}$ (38)
The values of these corner functions are equal to the values obtained from LNs, equations (35) and (36), within a range of a few times the respective SDs. However, the ratios of $\tilde{b}(3\pi/4)$ to the values obtained from LN are 0.86 or 0.6, which is not a sufficient agreement, and as the relative standard deviations are as large as 0.23 or 0.34, they are not stable values.
At the end of the calculation for a lattice satisfying PBCs, we take a square region $D$ whose centroid is $(99.5,99.5)$ with the site at the lower left corner of $U$ as the origin and repeat the same calculations as above. The results showed that for all corner functions treated, the values and their SDs are in agreement with (35), (36), (37) and (38) up to the most significant digit of each SD that are to be expected with the universal system $U$ under PBCs.
3 Harmonic lattice with fixed boundary conditions
In this section, as an universal system $U$, we consider a set of sites in a square satisfying FBCs of the same scale as in the previous section. Although the square has 1202 sites on a side, due to the FBCs for the two sites at each end, the number of sites with dynamic variables is 1200 and $L_{1},L_{2}=1201$. Also, for the subsystems $A_{1}$, $_{2}, A=A$$A_{1}+A_{2}$, we have the same shapes and scales as in the previous section (for $A_{2}$, square $A_{s}$, right isosceles triangle $A_{r}$, trapezoid $A_{t}$, parallelogram $A_{p}$). On the other hand, the region $D$, which includes $A$ and for which the correlation functions of Eqs.(19) and (20) are to be calculated is taken at several locations from the centroid of $U$ to near the lower left corner of $U$. We adopt the six positions of $D$ in Fig.2 as well as two regions whose centroid are (55.5,55.5) and (75.5,75.5) with the origin at the fixed end site at the lower left corner of $U$. The last two positions of $D$ are added to observe the fixed end effect to corner functions. For each of them, we compute $b_{\text{total}}$s and $\tilde{b}_{\text{total}}$s in Eqs.(31),(32) with figures of $A_{2}$ and corner functions involved as in the previous section.
@@ 3 Harmonic lattice with fixed boundary conditions
In the LN computations, for four $D$ regions closer to the centroid of $U$, the SDs are minimized when fitting to equation (32) up to $c_{-2}$ for $A_{s}$ and $A_{r}$ and up to $c_{-1}$ for $A_{t}$ and $A_{p}$, as in the case in where $U$ satisfies the PBCs. For the sake of consistent computation, we adopt as the value of each $b_{\text{total}}$ the value obtained by the same fitting as above for the four cases where the regions of $D$ are closer to the lower left corner of $U$. On the otherhand in the MI computations, for two $D$ regions closer to the centroid of $U$, the SDs are minimized when fitting up to $c_{-1}$ for all shapes of subsystem $A_{2}$ and so, we adopt as $\tilde{b}_{\text{total}}$ the values fitting up to $c_{-1}$ for all shapes of the other six $D$ regions, too. Each graph in Fig.3 plots the variation of each corner function corresponding to the location of the $D$ region with each SD as its error bar.
The values of corner functions $\pm$ their SDs obtained from LNs calculated in the D region with the cenroid $(600.5,600.5)$ are as follows:
$\begin{split}b(\pi/2)=0.02955\ \pm\ 0.00011,\\ b(\pi/4)=0.09670\ \pm\ 0.00032,\end{split}$ (39)
and as for two values obtained from $A_{t}$ and $A_{p}$,
$b(3\pi/4)=0.00604\pm 0.00047,\ \ b(3\pi/4)=0.00635\pm 0.00038..$ (40)
On the other hand, those values obtained from MIs are given by:
$\begin{split}\tilde{b}(\pi/2)=0.02963\pm 0.00003,\\ \tilde{b}(\pi/4)=0.1004\pm 0.0011,\end{split}$ (41)
and the two values of $\tilde{b}(3\pi/4)$ from $A_{t}$, $A_{p}$ are given by
$\tilde{b}(3\pi/4)=0.0052\pm 0.0013,\ \ \tilde{b}(3\pi/4)=0.0038\pm 0.0013.$ (42)
These values, respectively, are almost the same as those in Eqs.(35) $\sim$ Eq.(38) ( i.e., the values obtained in the universal system with PBCs and the $D$ region with the centroid (600.5,600.5). In each graph of Fig.3, each corner function obtained in the universal system with PBCs is also plotted at the horizontal axis value of 600, but it is indistinguishable from the plotted values of Eq(39) $\sim$ Eq(42), because they almost overlap. The results are expected, since, even when the universal system $U$ follows the FBCs, the effects of the fixed ends are expected to be small if the $D$ is sufficiently far from the boundary of the $U$. Furthermore, the values of each corner function within the horizontal axis coordinates (i.e., the center of the D region) between 300 and 600 are equal within their respective standard deviations, where the graph forms a horizontal line segment. Thus, for a sufficiently large system $U$, the effects of the fixed edge are to decrease in the region away from the edge of the $U$ about three to four times the size of the largest subsystem $A$, and the physical states of subsystems are expected to be the same as those of the harmonic lattice in the universal system obeying the PBCs.
In the conformal theories(CFTs), i.e. , the continuous limit of the above lattice theories, each corner function $\tilde{b}(\theta)$ by Renyi entropy with the order $1/2$ has a universal lower limit, $\tilde{b}\geq\frac{1}{32\pi}(\pi-\theta)^{2}$Bueno . Specifically,
Looking at the plots with the horizontal axis coordinate 600 of each graph in Fig.3, the error bars, i.e., SDs of $b(\pi/2)$, $\tilde{b}(\pi/2)$ and $b(3\pi/4)$, $\tilde{b}(3\pi/4)$ calculated from trapezoidal $A_{2}$ in shape, $A_{t}$, overlap respectively, where the lower limits expected by the CFT are included. On the other hand, for corner functions of $\pi/4$ and those of $3\pi/4$ calculated from parallelograms, $A_{p}$, when the error bars are set to three or two times SDs, where includes the lower limit expected by the CFT, they overlap.
Figure 3: Values of corner functions obtained in various $D$ regions. A scale value of the horizontal axis of each graph indicates (the $x$-coordinate of the centroid of $D$ region)0.5-0.5. The graphs in the lower row are values of $b(3\pi/4)$ obtained using $A_{t}$ and $A_{p}$, respectively.
4 3D lattice systems: Edge contributions to entanglement
So far, taking the universal system $U$ and the subsystems $A=A_{1}+A_{2}$ on a 2D lattice, we have analyzed the LNs and the MIs between $A_{1}$ and $A_{2}$. In this section, we investigate the entanglement between two adjacent sets of harmonic oscillators placed on lattice sites in $d_{U}=3$ dimensional space. For solid figures, the number of sites in the interior increases rapidly in proportion to the cube of the scale. Considering the limitation of available computer resources, we assume that the universal system $U$ is the set of sites in a cube with 120 sites on a edge. As a subsystem, we take the sets of sites near the center of gravity of the system $U$ that are inside a cube or a triangular prism whose base is a isosceles right triangle. Then, assuming $U=A+B$, we compute the Renyi entropy of order $1/2$ while varying the scale,the number of sites on the shortest edge of each shape of the subsystems $A$ from 6 to 20. In these cases, the boundary surfaces of the two subsystems $A$ and $B$ have not only sharp vertices but also edges. As for the contribution to entanglement from sharp corners and edges, the following equation is said to be valid from formula (6) omitting cut-offDevakul ; Sierens :
$S_{A}^{(1/2)}=\alpha P_{B}+\beta_{total}\log P_{B}-\sum_{j}e(\theta_{j})E_{j}+\cdots,$ (43)
where $\beta_{total}$ is the sum of corner function $\beta(\Omega_{k})$ of sharp vertex with the solid angle $\Omega_{k}$, and $e(\theta_{j})$ and $E_{j}$ are the edge term coefficient and the length of the edges with dihedral angle $\theta_{j}$. By fitting the above data of the Renyi entropy according to the formula (43), we obtain the corner functions $\beta(\pi/2)$, $\beta(\pi/4)$, the edge term coefficients $e(\pi/2)$ and $e(\pi/4)$ and their SDs. When doing so, as is done in equations (31), (32) and later, terms proportional to $P_{B}^{-k}(k=0,1,2...,n)$ are added to the right-hand side of the equation (43), and the operations are repeated while increasing $n$. Then, we select the $n$ that gives the most stable total values of corner functions, i.e., the smallest SD of the total value $\beta_{total}$. In fact, in both the PBC and FBC cases and for both $A$ shapes of cube and triangle prism, the most stable total corner functions are obtained by the fitting up to the constant term $c_{0}$.
To begin, for the case where the shape of the subsystem $A$ is a cube, let us compare the sum of the corner functions and their SDs obtained by the method described above with those obtained by fitting to the Eq.(43) minus the edge terms. The table 1 gives the corner functions and their SDs when computing with and without edge terms for the case where $U$ satisfies the PBCs and FBCs, respectively. In the case where the edge terms are considered, the coefficient values and their SDs are also included.
$\beta(\pi/2)\pm\beta(\pi/2)_{sd}$ $e(\pi/2)\pm e(\pi/2)_{sd}$
PBC without $e$ $0.242\pm 0.026$
with $e$ $005673\pm 0.00010 0.0$$06725\pm 0.00003$
FBC without $e$ $0.242\pm 0.026$
with $e$ $005604\pm 0.000059 0.0$6723±0.000020.06723\ $\pm$ 0.00002
Although the corresponding values in the PBC row and FBC row of the table are almost the same, the corner functions without considering the edge terms are about 43 times larger than those with the edge terms. Furthermore the relative SDs of corner functions are about 0.11 without considering the edge terms, while those are about 0.02 for PBCs and 0.01 for FBCs when the edge terms are considered. In other words, when the edge terms are taken into account, the corner functions are five to ten times more stable than when the edge terms are not taken into account. For the edge term coefficients, very stable values 0.06725, 0.06723 with relative SDs of 0.0004, 0.0003 are obtained for the two lattice boundary conditions. The results indicate the existence of the edge terms in the Eq.(43).
Next, let us discuss the case where the shape of $A$ system is a triangular prism. This triangular prism is with the base that is an isosceles right triangle formed by dividing one face (square) of the cube described above by its diagonal, and the height being the length of one edge of the cube. This shape has four and two sharp vertices with solid angle $\pi/4$, $\pi/2$, respectively, and two and seven edges with dihedral angle $\pi/4$, $\pi/2$, respectively. For the system of this configuration, we follow the same process as before, giving the sum of the related corner functions and edge terms, and using values of $\beta(\pi/2)$ and $e(\pi/2)$ and their SDs in the table 1, we obtain corner functions $\beta(\pi/4)$ and edge term coefficients $e(\pi/4)$ and their SDs as shown in the table 2.
$\beta(\pi/4)\pm\beta(\pi/4)_{sd}$ $e(\pi/4)\pm e(\pi/4)_{sd}$
PBC $0.0115\pm 0.0002$ $0.1416\pm 0.0001$
FBC $0.0114\pm 0.0001$ $0.1417\pm 0.0001$
The values of corner functions, the edge term coefficients are consistent within the range of their SDs between two cases where the universal lattice satisfies PBCs and FBCs, and both of them are stable values with relative SDs below 0.02 or 0.001. In addition, the values of $b(\pi/4)$ and $e(\pi/4)$ are about twice as large as those of $b(\pi/2)$ and $e(\pi/2)$, respectively. These characteristics are consistent with the fact that the geometric discontinuity of the boundary surface becomes stronger as the solid angle or dihedral angle approaches an acute angle, and the contribution from sharp corners and edges increases, as was the case with the two dimensional lattice.
5 conclusions
We have studied the logarithmic negativity(LN) and mutual information(MI) between two adjacent subsets $A_{1}$ and $A_{2}$ in a isolated universal set $U$ of harmonic oscillators on a two dimensional lattice, whose continuum limit is the massless scalar field. We have let the sites of $A_{2}$ be enclosed by sites of $A_{1}$ and let the boundary of $A_{2}$ be one of various shapes with corners whose angles are $\pi/4$, $\pi/2$, $3\pi/4$. Varying the scales of subsystems $A=A_{1}+A_{2}$ and $A_{2}$ kept as similar as possible in their shapes, we have obtained LNs, MIs and from them so-called corner functions for the above corner angles to verify the results of the previous study which treats corner functions mainly from LNs for the case where the canonical variables obey periodic boundary conditions(PBCs). In the case that canonical variables on the lattice obey PBCs, all the values of corner functions $b(\theta)$ obtained from LNs almost agree with those in the previous studies, although some of the differences are quite large relative to the standard deviations. As for the corner function values $\tilde{b}(\theta)$ obtained from MIs, the values of the corner function $\tilde{b}(\pi/2)$ and $\tilde{b}(\pi/4)$ are almost agreed with those obtained from data of $LN$s. On the other hand, as for the values of $\tilde{b}(3\pi/4)$, considering the magnitude of the SDs and their ratio to those obtained from $LN$s data, or the previous studyNobili , they are not stable and not enough consistent to those of the previous study.
@@
Furthermore, We have computed the Renyi entropies of order 1/2 for a three-dimensional lattice, considering, as a subsystem, a set of sites inside a cube varying the scale of the subsystem. By fitting the data with equation (43), we have obtained the corner functions and edge term coefficients as shown in table 1. We have compared the SDs of the corner functions with those obtained by fitting with (43) minus the edge terms . The result shows that the SDs of the latter are 5 to 10 times larger than that of the former. This fact is the evidence of the existence of the edge terms in the expansions of the Renyi entropy between two adjacent systems whose boundary is a three-dimensional polyhedron. In addition, the Renyi entropies also have been obtained for the sets of sites in triangular prisms with vertices and edges having solid angles and dihedral angles of $\pi/2$, $\pi/4$ as the subsystem $A$. Using the results of table 1, the corner function $\beta(\pi/4)$, the edge term coefficient $e(\pi/4)$ and their SDs have been obtained as shown in the table 2.
@@
As in the case of a two-dimensional lattice, the values of the corner functions and edge term coefficients in tables 1 and 2, where U satisfies FBCs and the $A$ system is sufficiently distant from the fixed ends of the universal system $U$, are almost identical to the values when $U$ satisfies PBCs. In addition, these values satisfy the stability and monotonically decreasing properties with respect to the solid angle and dihedral angle, respectively, that the corner functions and edge terms coefficients would have.